1 12. Special Journals are used to: a Increase efficiency b. Reduce costs c. Reduce the number of similar journal entries prepared d. All of the above e. None of the above 13. Segregation of duties is one example of a Gencrally Accepted Accounting Principles b. FASB c. Internal Controls d. Accounting Cycle 14. Which of the following is NOT a generally accepted accounting principle? a. Balance Principle b. Matching Principle c. Revenue recognition principle d. Consistency principle 15. When preparing bank reconciliation, you notice that a $500 check for rent cleared the bank correctly but was only listed on the books as S50. Which of the following is ncoded? a. Add to the bank balance b. Subtract from the bank balance c. Add to the book balance d. Subtract from the book balance 16. Howc Company has the following inventory transactions during the month. Beginning Inventory 05/15 Purchase 05/20 Sold 05/22 Purchase 05/30 Sold 50 units @ $10-unit cost 100 units $12-unit cost 60 units 0 units&$13-unit cos 80 units 17. What is the valuc of the ending inventory using First in First out for a periodic Inventory System? a. $620 b. $770 С. $775 d. $780 3
Answer #1

Question 12

Correct answer is option (d): All of the above

Special journals are standard templates prepared for frequently occurring transactions. By doing so, the need to create a new journal is eliminated, every time that frequent transaction occurs. This leads to speed and efficiency improvement. This also prevents the need to prepare new journal and thus reduce costs. Hence all the benefits stated in the options are achieved by special journal.

Question 13

Correct answer is option (c): Internal control

Segregation of duties (SOD) is an internal control that requires different people to complete different parts of a task. It is designed to ensure that at least two individuals are responsible for the separate parts of any task. It involves breaking down tasks that might reasonably be completed by a single individual into multiple tasks so that no one person is solely in control. It tries to ensure that a person is not in a position to perpetrate and conceal the errors or fraud.

Question 14

Correct answer is option (a): balance principle

Matching principle requires matching the revenues and expenses of the period so that profits for a period are stated properly on the financial statements. Revenue recognition principle states recognition of revenues only when certain criteria are met. Consistency principle requires a consistency in reporting format. If certain changes are required, prior period figures need to be restated to bring consistency in the reporting. There is nothing called balance principle hence, it's not a generally acceptable accounting principle.

Question 15

Correct answer is option (d): Subtract from the book balance.

There is no error from bank side. Hence bank balance will already be correct. The error is in the book balance and the same should be rectified by subtracting the shortfall accounted amount = $ 500 - $ 50 = $ 450 from the book balance so that bank balance and book balance match.

Question 16

Correct answer is option (b): $ 770

60 units sold on 05/20 should first be accounted against  opening inventory of 50 units and then balance 10 units towards units purchased on 05/15. So, we will now 90 units purchased on 05/15 balance in inventory. 80 units sold on 05/30 should be first accounted against the units balance from purchases on 05/15. So, inventory balance at the end of the period =

  • 10 units from purchases on 05/15; valuation = 10 units x $ 12 per unit = $ 120
  • All the units purchased on 05/22; valuation = 50 units x $ 13 per unit = $ 650
  • Hence total inventory value at close = 120 + 650 = $ 770